# Emergent chirality and superfluidity of parity-doubled baryons in neutron stars

### v1

@@ II.2 Emergent chiral symmetry
$\psi_{\mathrm{L}}\rightarrow e^{i\tau_{3}\theta_{\mathrm{L}}}\psi_{\mathrm{L}},\quad\psi_{\mathrm{R}}\rightarrow e^{i\tau_{3}\theta_{\mathrm{R}}}\psi_{\mathrm{L}},\quad$ ${\rm with}\quad(e^{i\tau_{3}\theta_{\mathrm{L}}},e^{i\tau_{3}\theta_{\mathrm{R}}})\in\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}},$ (2)
@@
for each chirality. The discrete group ${\mathbb{Z}}_{2\alpha}^{\prime}$ in the denominators on the right hand sides is necessary, because a simultaneous action of the $\pi$ rotation $\psi_{\alpha}\rightarrow e^{i\tau_{3}\pi}\psi_{\alpha}=-\psi_{\alpha}$ of $\mathrm{U}(1)_{(1-2)\alpha}$ and that $e^{i\pi}\psi_{\alpha}=-\psi_{\alpha}$ of $\mathrm{U}(1)_{\alpha}$ is redundant, and should be removed. In this paper, we shall name the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ symmetry an emergent chiral symmetry.
The ordinary chiral symmetry $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}$ is of course a subgroup of the total global symmetry by definition:
@@
The $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry expresses the internal symmetry for the doublet $(\psi_{1},\psi_{2})^{\mathrm{t}}$. This is a novel symmetry which has not been recognized in the past studies [12, 13]. The $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry reproduces the naive assignment and the mirror assignment as the subgroups when a certain symmetry-locking is introduced: if our theory respects only a simultaneous phase rotation of $\mathrm{U}(1)_{1\mathrm{L}}$ and $\mathrm{U}(1)_{2\mathrm{L}}$ and that of $\mathrm{U}(1)_{1\mathrm{R}}$ and $\mathrm{U}(1)_{2\mathrm{R}}$, it gives the naive assignment, while if it does only that of $\mathrm{U}(1)_{1\mathrm{L}}$ and $\mathrm{U}(1)_{2\mathrm{R}}$ and that of $\mathrm{U}(1)_{2\mathrm{L}}$ and $\mathrm{U}(1)_{1\mathrm{R}}$, it gives the mirror assignment. Therefore, the naive and mirror assignments can be identified as the subgroups $\mathrm{U}(1)_{1\mathrm{L}+2\mathrm{L}}\times\mathrm{U}(1)_{1\mathrm{R}+2\mathrm{R}}$ and $\mathrm{U}(1)_{1\mathrm{L}+2\mathrm{R}}\times\mathrm{U}(1)_{1\mathrm{R}+2\mathrm{L}}$, respectively. In this sense, the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry will provide us with a unified picture for the parity-doubled baryons. Given that the mirror assignment is favored than the naive assignment in the recent lattice QCD simulations at finite temperature [14, 15, 16] [^3] As a related work, the parity-doublet model is adopted also for the fluctuations around the critical point in the QCD phase diagram [80]., we may naturally ask how the emergent chiral symmetry, as an extension of the usual chiral symmetry, can appear in a variety of QCD phases at finite temperature, finite density, external fields and so on. To this end, we shall study how the emergent chiral symmetry arises in various phenomena of hadronic matter.
@@
Here ${\cal L}_{\mathrm{com}}$ is the one including only the terms invariant commonly in the naive and mirror assignments. In contrast, $\delta{\cal L}_{\mathrm{naive}}$ ($\delta{\cal L}_{\mathrm{mirror}}$) is the term invariant for the naive (mirror) assignment as a subgroup of the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry, but it composed of only the terms which are not included in ${\cal L}_{\mathrm{com}}$. The Lagrangians invariant in the naive and mirror assignments are given by ${\cal L}_{\mathrm{naive}}={\cal L}_{\mathrm{com}}+\delta{\cal L}_{\mathrm{naive}}$ and ${\cal L}_{\mathrm{mirror}}={\cal L}_{\mathrm{com}}+\delta{\cal L}_{\mathrm{mirror}}$, respectively. See also Fig. 2. Note that $\delta{\cal L}_{\mathrm{naive}}$ and $\delta{\cal L}_{\mathrm{mirror}}$ break the symmetry of ${\cal L}_{\mathrm{com}}$ explicitly. In the present study, we will focus on the ${\cal L}_{\mathrm{com}}$ to clarify generic properties shared by the naive and mirror assignments as a first step of our study.
@@
To describe the emergent chiral superfluids, we introduce the four-point interaction for $\psi_{1}$ and $\psi_{2}$. Thus we consider the Lagrangian invariant under the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry:
@@
The invariance of ${\cal L}_{\mathrm{com}}$ holds also for both of the naive and mirror assignments. Here we have three interaction terms with the coupling constants, $g_{\perp}$, $g_{\perp}^{\prime}$ and $g_{\parallel}$, which take different values in general. The coefficients are multiplied by constant numbers, $-4$ and $-8$, for the later convenience. Similarly, we can construct the Lagrangians ${\cal L}_{\mathrm{naive}}$ and ${\cal L}_{\mathrm{mirror}}$ for the naive and mirror assignments, respectively, as shown in Eqs. (61) and (62) in the Appendix B. In this study, we rather concentrate on the ${\cal L}_{\mathrm{com}}$ to clarify the common properties shared by the naive and mirror assignments.
Besides the $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}$ chiral symmetry, it is natural to introduce a further internal symmetry for the two fermions $\psi_{1}$ and $\psi_{2}$ to impose additional constraints in the Lagrangian (7). The requirement of such an internal symmetry stems from the observation that there is no a priori criterion to distinguish $\psi_{1}$ and $\psi_{2}$ in essence, and hence some internal symmetry may naturally be arisen to exchange $\psi_{1}$ and $\psi_{2}$. Referring the decompositions in Eqs. (4), the above consideration leads to enlargement of the emergent chiral symmetry, $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$, to higher symmetry to realize the exchange of $\psi_{1}$ and $\psi_{2}$. The higher emergent chiral symmetry represents the additional constraint for the exchange of $\psi_{1}$ and $\psi_{2}$, and reduces the number of terms in the Lagrangian (7).
@@
with an $\mathrm{SU}(2)$ symmetry being continuously exchanging $\psi_{1}$ and $\psi_{2}$ for each chirality: $\psi_{\alpha}\rightarrow e^{i\boldsymbol{\tau}\cdot\boldsymbol{\theta}_{\alpha}/2}\psi_{\alpha}$ ($\alpha=\mathrm{L},\mathrm{R}$) with the Pauli matrices, $\boldsymbol{\tau}=(\tau_{1},\tau_{2},\tau_{3})$, and the real parameters, $\boldsymbol{\theta}_{\alpha}=(\theta_{1\alpha},\theta_{2\alpha},\theta_{3\alpha})$.
Since the emergent chiral symmetries $G_{\mathrm{L}}\times G_{\mathrm{R}}$ are extensions of $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$, their actions have a redundancy with that of the chiral symmetry $\mathrm{U}(1)_{1\mathrm{R}}\times\mathrm{U}(1)_{2\mathrm{R}}$ that should be removed as in Eq. (5):
@@
where the original chiral symmetry $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}$ in the first line is extended into the $G_{\mathrm{L}}\times G_{\mathrm{R}}=\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetries (see Table 1 and Fig. 3). In the followings, we study the consequences of the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetries and their thermodynamics.
@@
Table 1: The full chiral symmetries are summarized: $\frac{\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{L}}}{{\mathbb{Z}}_{2\mathrm{L}}^{\prime}}\times\frac{\mathrm{U}(1)_{\mathrm{R}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}}{{\mathbb{Z}}_{2\mathrm{R}}^{\prime}}$ and $\mathrm{U}(2)_{\mathrm{L}}\times\mathrm{U}(2)_{\mathrm{R}}$ for the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetries, respectively, where the actions of their universal covering groups $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}\times\mathrm{SU}(2)_{\mathrm{R}}$ are shown. $\psi_{\alpha}^{\mathrm{t}}=(\psi_{1\alpha},\psi_{2\alpha})^{\mathrm{t}}$ is the doublet ($\alpha=\mathrm{L},\mathrm{R}$). $\boldsymbol{\tau}=(\tau_{1},\tau_{2},\tau_{3})$ are the Pauli matrices in the $\mathrm{SU}(2)$ space, $\varphi_{\alpha}$ is a real number and $\boldsymbol{\theta}_{\alpha}=(\theta_{1\alpha},\theta_{2\alpha},\theta_{3\alpha})$ are real parameters.
Figure 3: The inclusion relations are shown for the Lagrangians with $\frac{\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{L}}}{{\mathbb{Z}}_{2\mathrm{L}}^{\prime}}\times\frac{\mathrm{U}(1)_{\mathrm{R}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}}{{\mathbb{Z}}_{2\mathrm{R}}^{\prime}}$ for the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry, and $\mathrm{U}(2)_{\mathrm{L}}\times\mathrm{U}(2)_{\mathrm{R}}$ for the $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetry. In the present study, we consider the Lagrangian terms commonly shared by the naive and mirror assignments as indicated by solid lines (com). See also Fig. 2. $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}$ indicates $\mathrm{U}(1)_{1\mathrm{L}+2\mathrm{L}}\times\mathrm{U}(1)_{1\mathrm{R}+2\mathrm{R}}$ in the naive assignment and $\mathrm{U}(1)_{1\mathrm{L}+2\mathrm{R}}\times\mathrm{U}(1)_{1\mathrm{R}+2\mathrm{L}}$ in the mirror assignment. The $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{1\mathrm{R}}$ symmetric term included commonly in both naive and mirror assignments, but possessing no $\frac{\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{L}}}{{\mathbb{Z}}_{2\mathrm{L}}^{\prime}}\times\frac{\mathrm{U}(1)_{\mathrm{R}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}}{{\mathbb{Z}}_{2\mathrm{R}}^{\prime}}$ symmetry, is shown by the $g_{1212}$ term in Eqs. (61) and (62) in Appendix B.
II.3 Lagrangisns with the emergent chiral symmetries

### v2

@@ II.2 Emergent chiral symmetry
$\psi_{\mathrm{L}}\rightarrow e^{i\tau_{3}\theta_{\mathrm{L}}}\psi_{\mathrm{L}},\quad\psi_{\mathrm{R}}\rightarrow e^{i\tau_{3}\theta_{\mathrm{R}}}\psi_{\mathrm{R}},\quad$ ${\rm with}\quad(e^{i\tau_{3}\theta_{\mathrm{L}}},e^{i\tau_{3}\theta_{\mathrm{R}}})\in\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}},$ (2)
@@
for each chirality. The discrete group ${\mathbb{Z}}_{2\alpha}^{\prime}$ in the denominators on the right-hand sides is necessary, because a simultaneous action of the $\pi$ rotation $\psi_{\alpha}\rightarrow e^{i\tau_{3}\pi}\psi_{\alpha}=-\psi_{\alpha}$ of $\mathrm{U}(1)_{(1-2)\alpha}$ and that $e^{i\pi}\psi_{\alpha}=-\psi_{\alpha}$ of $\mathrm{U}(1)_{\alpha}$ is redundant and should be removed. In this paper, we shall name the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ symmetry an emergent chiral symmetry.
The ordinary chiral symmetry $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}$ is of course a subgroup of the total global symmetry by definition
@@
The $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry expresses the internal symmetry for the doublet $(\psi_{1},\psi_{2})^{\mathrm{t}}$. This is a novel symmetry which has not been recognized in the past studies [12, 13]. The $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry reproduces the naive assignment and the mirror assignment as the subgroups when a certain symmetry-locking is introduced: if our theory respects only a simultaneous phase rotation of $\mathrm{U}(1)_{1\mathrm{L}}$ and $\mathrm{U}(1)_{2\mathrm{L}}$ and that of $\mathrm{U}(1)_{1\mathrm{R}}$ and $\mathrm{U}(1)_{2\mathrm{R}}$ it gives the naive assignment, while if it does only that of $\mathrm{U}(1)_{1\mathrm{L}}$ and $\mathrm{U}(1)_{2\mathrm{R}}$ and that of $\mathrm{U}(1)_{2\mathrm{L}}$ and $\mathrm{U}(1)_{1\mathrm{R}}$, it gives the mirror assignment. Therefore, the naive and mirror assignments can be identified as the subgroups $\mathrm{U}(1)_{1\mathrm{L}+2\mathrm{L}}\times\mathrm{U}(1)_{1\mathrm{R}+2\mathrm{R}}$ and $\mathrm{U}(1)_{1\mathrm{L}+2\mathrm{R}}\times\mathrm{U}(1)_{1\mathrm{R}+2\mathrm{L}}$, respectively. In this sense, the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry will provide us with a unified picture for the parity-doubled baryons. Given that the mirror assignment is more favored than the naive assignment in the recent lattice QCD simulations at finite temperature [14, 15, 16]  [^5] As a related work, the parity-doublet model is adopted also for the fluctuations around the critical point in the QCD phase diagram [83]., we may naturally ask how the emergent chiral symmetry, as an extension of the usual chiral symmetry, can appear in a variety of QCD phases at finite temperature, finite density, external fields, and so on. To this end, we shall study how the emergent chiral symmetry arises in various phenomena of hadronic matter.
@@
Here ${\cal L}_{\mathrm{com}}$ is the one including only the terms invariant commonly in the naive and mirror assignments. In contrast, $\delta{\cal L}_{\mathrm{naive}}$ ($\delta{\cal L}_{\mathrm{mirror}}$) is the term invariant for the naive (mirror) assignment as a subgroup of the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry, but it is composed of only the terms which are not included in ${\cal L}_{\mathrm{com}}$. The Lagrangians invariant in the naive and mirror assignments are given by ${\cal L}_{\mathrm{naive}}={\cal L}_{\mathrm{com}}+\delta{\cal L}_{\mathrm{naive}}$ and ${\cal L}_{\mathrm{mirror}}={\cal L}_{\mathrm{com}}+\delta{\cal L}_{\mathrm{mirror}}$, respectively. See also Fig. 2. Note that $\delta{\cal L}_{\mathrm{naive}}$ and $\delta{\cal L}_{\mathrm{mirror}}$ break the symmetry of ${\cal L}_{\mathrm{com}}$ explicitly. In the present study, we will focus on the ${\cal L}_{\mathrm{com}}$ to clarify generic properties shared by the naive and mirror assignments as a first step of our study.
@@
To describe the emergent chiral superfluids, we introduce the four-point interaction for $\psi_{1}$ and $\psi_{2}$. Thus we consider the Lagrangian invariant under the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry
@@
The invariance of ${\cal L}_{\mathrm{com}}$ holds also for both of the naive and mirror assignments. Here we have three interaction terms with the coupling constants, $g_{\perp}$, $g_{\perp}^{\prime}$, and $g_{\parallel}$, which take different values in general. The coefficients are multiplied by constant numbers, $-4$ and $-8$, for the later convenience. Similarly, we can construct the Lagrangians ${\cal L}_{\mathrm{naive}}$ and ${\cal L}_{\mathrm{mirror}}$ for the naive and mirror assignments, respectively, as shown in Eqs. (61) and (62) in Appendix B. In this study, we rather concentrate on the ${\cal L}_{\mathrm{com}}$ to clarify the common properties shared by the naive and mirror assignments.
Besides the $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}$ chiral symmetry, it is natural to introduce a further internal symmetry for the two fermions $\psi_{1}$ and $\psi_{2}$ to impose additional constraints in the Lagrangian (7). The requirement of such an internal symmetry stems from the observation that there is no a priori criterion to distinguish $\psi_{1}$ and $\psi_{2}$ in essence, and hence some internal symmetry may naturally arise to exchange $\psi_{1}$ and $\psi_{2}$. Referring to the decompositions in Eq. (4), the above consideration leads to enlargement of the emergent chiral symmetry, $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$, to higher symmetry to realize the exchange of $\psi_{1}$ and $\psi_{2}$. The higher emergent chiral symmetry represents the additional constraint for the exchange of $\psi_{1}$ and $\psi_{2}$ reduces the number of terms in the Lagrangian (7).
@@
with an $\mathrm{SU}(2)$ symmetry continuously exchanging $\psi_{1}$ and $\psi_{2}$ for each chirality: $\psi_{\alpha}\rightarrow e^{i\boldsymbol{\tau}\cdot\boldsymbol{\theta}_{\alpha}/2}\psi_{\alpha}$ ($\alpha=\mathrm{L},\mathrm{R}$) with the Pauli matrices, $\boldsymbol{\tau}=(\tau_{1},\tau_{2},\tau_{3})$, and the real parameters, $\boldsymbol{\theta}_{\alpha}=(\theta_{1\alpha},\theta_{2\alpha},\theta_{3\alpha})$.
Since the emergent chiral symmetries $G_{\mathrm{L}}\times G_{\mathrm{R}}$ are extensions of $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$, their actions have a redundancy with that of the chiral symmetry $\mathrm{U}(1)_{1\mathrm{R}}\times\mathrm{U}(1)_{2\mathrm{R}}$ that should be removed as in Eq. (5)
@@
where the original chiral symmetry $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}$ in the first line is extended into the $G_{\mathrm{L}}\times G_{\mathrm{R}}=\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetries (see Table 1 and Fig. 3). In the following, we study the consequences of the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetries and their thermodynamics.
@@
Table 1: Full chiral symmetries are summarized: $\frac{\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{L}}}{{\mathbb{Z}}_{2\mathrm{L}}^{\prime}}\times\frac{\mathrm{U}(1)_{\mathrm{R}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}}{{\mathbb{Z}}_{2\mathrm{R}}^{\prime}}$ and $\mathrm{U}(2)_{\mathrm{L}}\times\mathrm{U}(2)_{\mathrm{R}}$ for the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetries, respectively, where the actions of their universal covering groups $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}\times\mathrm{SU}(2)_{\mathrm{R}}$ are shown. $\psi_{\alpha}^{\mathrm{t}}=(\psi_{1\alpha},\psi_{2\alpha})^{\mathrm{t}}$ is the doublet ($\alpha=\mathrm{L},\mathrm{R}$). $\boldsymbol{\tau}=(\tau_{1},\tau_{2},\tau_{3})$ are the Pauli matrices in the $\mathrm{SU}(2)$ space, $\varphi_{\alpha}$ is a real number, and $\boldsymbol{\theta}_{\alpha}=(\theta_{1\alpha},\theta_{2\alpha},\theta_{3\alpha})$ are real parameters.
Figure 3: Inclusion relations are shown for the Lagrangians with $\frac{\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{L}}}{{\mathbb{Z}}_{2\mathrm{L}}^{\prime}}\times\frac{\mathrm{U}(1)_{\mathrm{R}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}}{{\mathbb{Z}}_{2\mathrm{R}}^{\prime}}$ for the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry and $\mathrm{U}(2)_{\mathrm{L}}\times\mathrm{U}(2)_{\mathrm{R}}$ for the $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetry. In the present study, we consider the Lagrangian terms commonly shared by the naive and mirror assignments as indicated by solid lines (com). See also Fig. 2. $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{\mathrm{R}}$ indicates $\mathrm{U}(1)_{1\mathrm{L}+2\mathrm{L}}\times\mathrm{U}(1)_{1\mathrm{R}+2\mathrm{R}}$ in the naive assignment and $\mathrm{U}(1)_{1\mathrm{L}+2\mathrm{R}}\times\mathrm{U}(1)_{1\mathrm{R}+2\mathrm{L}}$ in the mirror assignment. The $\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{1\mathrm{R}}$ symmetric term included commonly in both naive and mirror assignments but possessing no $\frac{\mathrm{U}(1)_{\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{L}}}{{\mathbb{Z}}_{2\mathrm{L}}^{\prime}}\times\frac{\mathrm{U}(1)_{\mathrm{R}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}}{{\mathbb{Z}}_{2\mathrm{R}}^{\prime}}$ symmetry is shown by the $g_{1212}$ term in Eqs. (61) and (62) in Appendix B.
II.3 Lagrangians with the emergent chiral symmetries